What is the frequency of green light, which has a wavelength of 4.9X10^2m
1
Expert's answer
2020-11-11T10:35:08-0500

Solution:

the frequency ƒ = the speed of light c / the wavelenght λ

the green light has a wavelight 4.9 ×10-7 m

ƒ = 3.0×108 m/s × 4.9×10-7 m = 6.12 ×1014 Hz

Answer: the frequency of green light has 6.12 ×1014 Hz